# thanks: Corresponding authorthanks: Corresponding author

### v1

@@ II.4 Gaussian superchannels
Recall that a superchannel is a completely positive linear map transforming any quantum channels into quantum channels; and a Gaussian superchannel is a superchannel transforming any Gaussian channels into Gaussian channels (X11 ). Write
${\mathcal{SGC}}(H)=\{\mbox{all Gaussian superchannels on }H\}.$
Xu X11 gave two different representations of Gaussian superchannels.
Theorem 1.
(X11 ) Any $N$-mode Gaussian superchannel $\Phi\in{\mathcal{SGC}}(H)$ can be represented by $\Phi\left({A,E,Y,\nu}\right)$, where $\nu\in{\mathbb{R}}^{2N}$, $A,E,Y\in\mathcal{M}_{2N}(\mathbb{R})$ satisfy $Y=Y^{\rm T}$, $EE^{\rm T}={I_{2N}}$ and
$Y+i\Omega_{N}-iA\Omega_{N}A^{\rm T}\geq 0,\ \ i\Omega_{N}-iE\Omega_{N}E^{\rm T}\geq 0.$ (10)
Moreover, for any $\phi\left({K,M,\mathbf{d}}\right)\in\mathcal{GC}(H)$, we have
(1) $\Phi\left({\phi\left({K,M,\mathbf{d}}\right)}\right)=\phi^{\prime}\left({K^{\prime},M^{\prime},\mathbf{d}^{\prime}}\right)$ with
$K^{\prime}=AK\Sigma_{N}E^{\rm T}\Sigma_{N},\ M^{\prime}=AMA^{\rm T}+Y,\ \mathbf{d}^{\prime}=A\mathbf{d}+\nu,$
[…]
(2) $\Phi\left(\phi\left({K,M,\mathbf{d}}\right)\right)={\phi_{2}}\circ\phi\circ{\phi_{1}}$ with some ${\phi_{1}}\left(K_{1},M_{1},\mathbf{d}_{1}\right)$ and ${\phi_{2}}\left(K_{2},M_{2},\mathbf{d}_{2}\right)\in{\mathcal{GC}}(H)$, where $K_{1}=\Sigma_{N}E^{\rm T}\Sigma_{N},$ $M_{1}=0,$ ${\mathbf{d}_{1}}=0,$ $K_{2}=A,$ $M_{2}=Y$, ${\mathbf{d}_{2}}=\upsilon.$
@@ III Gaussian steering-annihilating and Gaussian steering-breaking channels
Assume that $H,K$ are any infinite-dimensional separable complex Hilbert spaces. We say that any Gaussian channel $\phi\in{\mathcal{GC}}(H\otimes K)$ is Gaussian steering-annihilating if it sends all Gaussian states into Gaussian unsteerable states, that is, if $\phi(\mathcal{GS}(H\otimes K))\subseteq{\mathcal{GS}}_{\mathcal{US}}(H\otimes K)$; and any Gaussian channel $\psi\in\mathcal{GC}(H)$ is Gaussian steering-breaking if $(\psi\otimes I_{K})(\rho)$ is always unsteerable for all Gaussian states $\rho\in{\mathcal{GS}}(H\otimes K)$, that is, if $(\psi\otimes I_{K})(\mathcal{GS}(H\otimes K))\subseteq{\mathcal{GS}}_{\mathcal{US}}(H\otimes K)$.
Let
$\begin{array}[]{rl}&{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})\\ =\{&\mbox{all Gaussian steering-annihilating }\\ &\mbox{channels on }H_{A}\otimes H_{B}\}\end{array}$
and
$\begin{array}[]{rl}&{\mathcal{GC}}_{\mathcal{SB}}(H)\\ =&\{\mbox{ all Gaussian steering-breaking channels on }H\}.\end{array}$
By Definition 2, the following useful property is obvious.
@@
If $\phi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$, and $\psi\in{\mathcal{GC}}(H_{A}\otimes H_{B})$, then $\phi\circ\psi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$; if $\phi\in{\mathcal{GC}}_{\mathcal{SB}}(H)$ and $\psi\in{\mathcal{GC}}(H)$, then both $\phi\circ\psi$ and $\psi\circ\phi$ belong to ${\mathcal{GC}}_{\mathcal{SB}}(H)$.
@@
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}\left(H_{A}\otimes H_{B}\right)$ is any $(m+n)$-mode channel. If $\phi$ satisfies the condition
$M+({0_{2m}}\oplus i{\Omega_{n}})-iK\left(\Omega_{m}\oplus\Omega_{n}\right){K^{\rm T}}\geq 0,$ (11)
then $\phi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$.
@@
Assume that $\rho\in{\mathcal{GS}}\left(H_{A}\otimes H_{B}\right)$ is any Gaussian state with CM $\Gamma_{\rho}$ in Eq.(3). Then $\phi(\rho)$ has the CM $\Gamma_{\phi(\rho)}=K\Gamma_{\rho}{K^{\rm T}}+M$. As $\Gamma_{\rho}+i\left({{\Omega_{m}}\oplus{\Omega_{n}}}\right)\geq 0$, by the assumption (11), we have
@@
$\geq$ $K(-i(\Omega_{m}\oplus\Omega_{n}))K^{\rm T}+M+(0_{2m}\oplus i\Omega_{n})\geq 0.$
It follows that $\phi\left(\rho\right)\in{\mathcal{GS}}_{\mathcal{U}S}(H_{A}\otimes H_{B})$. So $\phi$ is Gaussian steering-annihilating. ∎
Notice that the condition (11) is only sufficient but not necessary for a channel being steering-annihilating. In fact, there exist Gaussian steering-annihilating channels which do not satisfy the condition (11).
@@
Then $\phi_{1}$ is steering-annihilating, but does not satisfy the condition (11).
@@
So $\phi$ does not satisfy the condition (11).
@@
which implies that $\phi$ is steering-annihilating.
While Ineq.(11) is not a necessary condition, the subsequent Ineq.(12) provides a sufficient and necessary condition for a Gaussian channel becoming steering-annihilating.
@@
Any $(m+n)$-mode Gaussian channel $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$ if and only if
@@
Proof.
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}\left(H_{A}\otimes H_{B}\right)$ is any Gaussian channel. If $\phi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$, then for any $\rho\in{\mathcal{GS}}\left(H_{A}\otimes H_{B}\right)$ with CM $\Gamma_{\rho}$, by Eqs.(5)-(6), one has
$K\Gamma_{\rho}K^{\rm T}+M\geq\pm i(0_{2m}\oplus\Omega_{n}),$
and so
holds for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$.
Note that $\Gamma_{\rho}\geq\pm i(\Omega_{m}\oplus\Omega_{n})$, and $\inf\limits_{\sigma\geq\pm i\Omega_{N}}{\bf w}^{\dagger}\sigma{\bf w}=|{\bf w}^{\dagger}\Omega_{N}{\bf w}|$ for all $N$-mode quantum states $\sigma$ PMGH . Thus, the above inequality implies
for all $\mathbf{w}\in\mathbb{C}^{2(m+n)},$ that is, Ineq.(12) holds.
On the other hand, if Ineq.(12) holds, then for any $\mathbf{w}\in\mathbb{C}^{2(m+n)}$, one gets
$\geq$ $|\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}|$
$\geq$ $\pm i\mathbf{w}^{{\dagger}}(0_{2m}\oplus\Omega_{n})\mathbf{w}.$
So $\Gamma_{\phi(\rho)}\geq\pm i(0_{2m}\oplus\Omega_{n})$, which means that $\phi(\rho)$ is unsteerable. It follows that $\phi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$. ∎
Note that
$\left|{{\varphi_{r}}}\right\rangle=\frac{1}{\cosh r}{\sum\limits_{j=0}^{\infty}{\left({\tanh r}\right)}^{j}}\left|j_{H}\right\rangle\left|j_{H}\right\rangle\in H\otimes H$
is a Gaussian squeezed pure state, where $\{|j_{H}\rangle\}$ is the Fock basis of $H$ and $r\in\mathbb{R}$ is the squeezed parameter. For any Gaussian channel $\phi\in{\mathcal{GC}}(H)$, define
${\rho_{\phi}}=(\phi\otimes I_{H})(\left|{{\varphi_{r}}}\right\rangle\left\langle{{\varphi_{r}}}\right|).$
Clearly, $\rho_{\phi}\in{\mathcal{GS}}(H\otimes H)$. By the unsteerability of $\rho_{\phi}$, we can give necessary and sufficient conditions for Gaussian steering-breaking channels.
@@
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}\left(H\right)$ is any N-mode Gaussian channel. Then the following statements are equivalent.
(1) $\phi\in{\mathcal{GC}}_{\mathcal{SB}}(H)$ is Gaussian steering-breaking.
(2) ${\rho_{\phi}}\in{\mathcal{GS}}_{\mathcal{US}}(H\otimes H)$.
@@
$M-iK\Omega_{N}K^{\rm T}\geq 0.$
Proof.
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}\left(H\right)$ is any $N$-mode Gaussian channel.
$\left(2\right)\Rightarrow\left(3\right)$: Assume that ${\rho_{\phi}}\in{\mathcal{GS}}_{\mathcal{US}}(H\otimes H)$. Note that the CM $\Gamma_{{\rho_{\phi}}}$ of $\rho_{\phi}$} has the form X11
[…]
@@
$\Gamma_{{\rho_{\phi}}}+(0_{2N}\oplus i{\Omega_{N}})\geq 0.$ (14)
It is well known that a Hermitian matrix $W=\left(\begin{array}[]{cc}W_{11}&W_{12}\\ W_{12}^{{\dagger}}&W_{22}\end{array}\right)\geq 0$ if and only if $W_{22}\geq 0$ and $W_{11}-W_{12}W_{22}^{-1}W_{12}^{{\dagger}}\geq 0$ (ZH ). Then Ineq.(14) implies
Note that $(T+cI)^{-1}=\frac{1}{c}(I-\frac{T}{c})+o(c^{-2})$ (S1 ), where $o(c^{-2})$ stands for the Landau little which will be neglected when taking to $c\to\infty$. So
[…]
$Y+i{\Omega_{N^{\prime}}}\geq 0$}
@@
$X+i{\Omega_{N}}-Z\left({Y+i{\Omega_{N^{\prime}}}}\right)^{-1}{Z^{\rm T}}\geq 0$}.
This implies
$KX{K^{\rm T}}+M-KZ{\left({Y+i{\Omega_{N^{\prime}}}}\right)^{-1}}{Z^{\rm T}}{K^{\rm T}}$
$\geq$ $K\left({-i{\Omega_{N}}}\right){K^{\rm T}}+M=M-iK{\Omega_{N}}{K^{\rm T}}\geq 0.$
Also note that
$KX{K^{\rm T}}+M\geq 0$
by the denifition of the channel $\phi$. It follows that ${\Gamma_{\left({\phi\otimes I_{H^{\prime}}}\right)(\rho)}}+i(0_{N}\oplus{\Omega_{N^{\prime}}})\geq 0$. So $(\phi\otimes{I_{H^{\prime}}})\left(\rho\right)$ is unsteerable, and hence $\phi$ is steering-breaking. ∎
In the end of this section, we will investigate the relationship between Gaussian steering-breaking channels, Gaussian steering-annihilating channels and maximal Gaussian unsteerable channels.
By their definitions, it is obvious that
${\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})\subset{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B}).$
As a consequence of Proposition 3, for any $\phi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$ and $\psi\in{\mathcal{GC}}_{\mathcal{SB}}(H_{A}\otimes H_{B})$, we have
$\phi\circ\psi\in{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})\cap{\mathcal{GC}}_{\mathcal{SB}}(H_{A}\otimes H_{B}),$
Next, take an $(m+n)$-mode constant channel $\Theta=\Theta(0,\Gamma_{0},\mathbf{d}_{0})$. Obviously, by Theorem 7, $\Theta$ is steering-breaking. However, if $\Gamma_{0}+(0_{2m}\oplus i\Omega_{n})\ngeq 0$, $\Theta$ is neither steering-annihilating nor maximal Gaussian unsteerable. So
${\mathcal{GC}}_{\mathcal{SB}}(H_{A}\otimes H_{B})\not\subset{\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})$
and
${\mathcal{GC}}_{\mathcal{SB}}(H_{A}\otimes H_{B})\not\subset{\mathcal{GC}}_{\mathcal{MUS}}(H_{A}\otimes H_{B}).$
@@
${\mathcal{GC}}_{\mathcal{SA}}(H_{A}\otimes H_{B})\not\subset{\mathcal{GC}}_{\mathcal{SB}}(H_{A}\otimes H_{B}).$
For example, take $\tilde{\phi}=\phi\otimes I_{B}\in{\mathcal{GC}}(H_{A}\otimes H_{B})$, where $\phi=\phi(\cos\theta I_{2},\sin^{2}\theta I_{2},\mathbf{d}_{0})$ is a single-mode attenuator channel. Then $\tilde{\phi}=\tilde{\phi}(K,M,\mathbf{d})$ can be represented by

### v2

@@ II.4 Gaussian superchannels
Recall that a superchannel is a completely positive linear map transforming any quantum channels into quantum channels; and a Gaussian superchannel is a superchannel transforming any Gaussian channels into Gaussian channels X11 .
Denote by
${\mathcal{GSC}}^{(N)}=\{\mbox{all N-mode Gaussian superchannels}\}.$
It is shown in X11 that any $N$-mode Gaussian superchannel $\Phi\in{\mathcal{GSC}}^{(N)}$ can be described by $\Phi\left({A,E,Y,\nu}\right)$ as follows: for any $N$-mode Gaussian channel $\phi\left({K,M,\mathbf{d}}\right)$, $\Phi\left({\phi\left({K,M,\mathbf{d}}\right)}\right)=\phi^{\prime}\left({K^{\prime},M^{\prime},\mathbf{d}^{\prime}}\right)$ with
$K^{\prime}=AK\Sigma_{N}E^{\rm T}\Sigma_{N},$
$M^{\prime}=AMA^{\rm T}+Y,$
$\mathbf{d}^{\prime}=A\mathbf{d}+\nu,$
where $A,E,Y\in\mathcal{M}_{2N}(\mathbb{R})$, $Y=Y^{\rm T}$, $EE^{\rm T}={I_{2N}}$, $\nu\in{\mathbb{R}}^{2N}$, $\Sigma_{N}=\underbrace{\Sigma\oplus\cdots\oplus\Sigma}_{N}$ with $\Sigma={\begin{pmatrix}1&0\\ 0&{-1}\end{pmatrix}}$, and
$Y+i\Omega_{N}-iA\Omega_{N}A^{\rm T}\geq 0,$
$i\Omega_{N}-iE\Omega_{N}E^{\rm T}\geq 0.$ (10)
In addition, the author X11 also proved that any $N$-mode Gaussian superchannel $\Phi\left({A,E,Y,\nu}\right)$ can be described in terms of compositions of Gaussian channels, that is, for any $N$-mode Gaussian channel $\phi\left({K,M,\mathbf{d}}\right)$, there exist two $N$-mode Gaussian channels ${\phi_{1}}\left(K_{1},M_{1},\mathbf{d}_{1}\right)$ and ${\phi_{2}}\left(K_{2},M_{2},\mathbf{d}_{2}\right)$ such that $\Phi(\phi)={\phi_{2}}\circ\phi\circ{\phi_{1}}$. Particularly, one such representation is
$\begin{cases}K_{1}=\Sigma_{N}E^{\rm T}\Sigma_{N},\ \ M_{1}=0,\ \ {\mathbf{d}_{1}}=0;\\ K_{2}=A,\ \ M_{2}=Y,\ \ {\mathbf{d}_{2}}=\upsilon.\end{cases}$
@@ III Gaussian steering-annihilating and Gaussian steering-breaking channels
Assume that $\phi$ is any $(m+n)$-mode bipartite Gaussian channel. We say that $\phi$ is steering-annihilating if it sends all $(m+n)$-mode Gaussian states into either $(m+n)$-mode Gaussian unsteerable states from A to B or $(m+n)$-mode Gaussian unsteerable states from B to A.
Definition 2.
Assume that $\psi$ is any $m$-mode Gaussian channel. We say that $\psi$ is steering-breaking if for any $(m+n)$-mode Gaussian state $\rho$, $(\psi\otimes I_{K})(\rho)$ is always $(m+n)$-mode Gaussian unsteerable from A to B or from B to A, where $K$ denotes arbitrary $n$-mode ancillary CV system.
For the convinience, denote by ${\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$, ${\mathcal{GC}}^{(m,n)}_{\mathcal{SA}(A\rightarrow B)}$ the set of all $(m+n)$-mode Gaussian steering-annihilating channels and the set of all $(m+n)$-mode Gaussian steering-annihilating channels sending all $(m+n)$-mode Gaussian states into $(m+n)$-mode Gaussian unsteerable states from A to B, respectively. Denote by ${\mathcal{GC}}^{(N)}_{\mathcal{SB}}$, ${\mathcal{GC}}^{(N)}_{\mathcal{SB}(A\rightarrow B)}$ the set of all $N$-mode Gaussian steering-breaking channels and the set of all $N$-mode Gaussian steering-breaking channels from A to B, respectively.
By Definitions 1 and 2, the following useful property is obvious.
@@
If $\phi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$, and $\psi\in{\mathcal{GC}}^{(m,n)}$, then $\phi\circ\psi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$; if $\phi\in{\mathcal{GC}}^{(N)}_{\mathcal{SB}}$ and $\psi\in{\mathcal{GC}}^{(N)}$, then both $\phi\circ\psi$ and $\psi\circ\phi$ belong to ${\mathcal{GC}}^{(N)}_{\mathcal{SB}}$.
@@
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}^{(m,n)}$ is any $(m+n)$-mode Gaussian channel. If $\phi$ satisfies one of the following conditions:
(1) $M+({0_{2m}}\oplus i{\Omega_{n}})-iK\left(\Omega_{m}\oplus\Omega_{n}\right){K^{\rm T}}\geq 0,$
(2) $M+(i{\Omega_{m}}\oplus{0_{2n}})-iK\left(\Omega_{m}\oplus\Omega_{n}\right){K^{\rm T}}\geq 0,$
then $\phi$ is steering-annihilating, that is, $\phi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$.
@@
Assume that $\rho\in{\mathcal{GS}}^{(m,n)}$ is any Gaussian state with CM $\Gamma_{\rho}$ in Eq.(3). Then $\phi(\rho)$ has the CM $\Gamma_{\phi(\rho)}=K\Gamma_{\rho}{K^{\rm T}}+M$. As $\Gamma_{\rho}+i(\Omega_{m}\oplus\Omega_{n})\geq 0$, we have
$K(\Gamma_{\rho}+i(\Omega_{m}\oplus\Omega_{n}))K^{\rm T}\geq 0$ (11)
for any $K\in\mathcal{M}_{2(m+n)}(\mathbb{R})$. If the condition (1) holds, combining this assumption with Ineq.(11), one gets
@@
$=$ $K(\Gamma_{\rho}+i(\Omega_{m}\oplus\Omega_{n}))K^{\rm T}$
$+(M+(0_{2m}\oplus i\Omega_{n})-K(i(\Omega_{m}\oplus\Omega_{n}))K^{\rm T})\geq 0.$
It follows from the steering criterion (Ineq.(5)) that $\phi\left(\rho\right)\in\mathcal{GS}^{(m,n)}_{\mathcal{US}(A\rightarrow B)}$.
Similarly, one can show that, if the condition (2) holds, then $\phi\left(\rho\right)\in\mathcal{GS}^{(m,n)}_{\mathcal{US}(B\rightarrow A)}$. Hence $\phi$ is Gaussian steering-annihilating. ∎
Notice that the conditions in Theorem 4 is only sufficient but not necessary for a channel being steering-annihilating. In fact, there exist Gaussian steering-annihilating channels which do not satisfy this condition.
@@
Then $\phi_{1}$ is steering-annihilating from A to B, but does not satisfy the condition (1) in Theorem 4.
@@
So $\phi$ does not satisfy the condition (1) in Theorem 4.
@@
which implies that $\phi$ is steering-annihilating (from A to B).
While the condition (1) or (2) in Theorem 4 is not a necessary condition, the subsequent inequality in Theorem 6 provides a sufficient and necessary condition for a Gaussian channel becoming steering-annihilating.
@@
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}^{(m,n)}$ is any $(m+n)$-mode Gaussian channel. Then $\phi$ is Gaussian steering-annihilating, that is, $\phi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$, if and only if
@@
holds for all $\mathbf{w}\in\mathbb{C}^{2(m+n)}$, or
$\geq$ $|\mathbf{w}^{{\dagger}}({\Omega_{m}}\oplus{0_{2n}})\mathbf{w}|$
@@
Take any $N$-mode Gaussian channel $\phi\in{\mathcal{GC}}^{(N)}$ with the state space $H$. Note that the pure state
$\left|{{\varphi_{r}}}\right\rangle=\frac{1}{\cosh r}{\sum\limits_{j=0}^{\infty}{\left({\tanh r}\right)}^{j}}\left|j\right\rangle\left|j\right\rangle$
is a two-mode Gaussian squeezed pure state, where $\{|j\rangle|j\rangle\}$ is a tensor product of Fock states and $r\in\mathbb{R}$ is the squeezed parameter. Correspondingly, $2N$-mode Gaussian squeezed pure state can be written as $\left|{{\psi_{r}}}\right\rangle=\left|{{\varphi_{r}}}\right\rangle^{\otimes N}.$ Define
${\rho_{\phi}}=(\phi\otimes I_{H})(\left|{{\psi_{r}}}\right\rangle\left\langle{{\psi_{r}}}\right|).$
Clearly, $\rho_{\phi}\in{\mathcal{GS}}^{(N,N)}$ is a $(N+N)$-mode Gaussian state. By the unsteerability of $\rho_{\phi}$, we can give necessary and sufficient conditions for Gaussian steering-breaking channels.
@@
Assume that $\phi=\phi\left({K,M,\mathbf{d}}\right)\in{\mathcal{GC}}^{(N)}$ is any $N$-mode Gaussian channel. Then the following statements are equivalent.
(1) $\phi$ is Gaussian steering-breaking, that is, $\phi\in{\mathcal{GC}}^{(N)}_{\mathcal{SB}}$.
(2) ${\rho_{\phi}}\in{\mathcal{GS}^{(N,N)}_{\mathcal{US}(A\rightarrow B)}}$ or ${\rho_{\phi}}\in{\mathcal{GS}^{(N,N)}_{\mathcal{US}(B\rightarrow A)}}$.
@@
$M-iK\Omega_{N}K^{\rm T}\geq 0$
or
$M+i\Omega_{N}\geq 0.$
Proofs of Theorems 6 and 7 are given in Appendix.
In the end of this section, we will investigate the relationship between Gaussian steering-breaking channels, Gaussian steering-annihilating channels and maximal Gaussian unsteerable channels. Denote by
${\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}}={\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(A\rightarrow B)}\bigcup{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}(B\rightarrow A)}.$
By their definitions, it is obvious that
${\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}={\mathcal{GC}}^{(m,n)}_{\mathcal{SA}(A\rightarrow B)}\bigcup{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}(B\rightarrow A)},$
${\mathcal{GC}}^{(m,n)}_{\mathcal{SB}}={\mathcal{GC}}^{(m,n)}_{\mathcal{SB}(A\rightarrow B)}\bigcup{\mathcal{GC}}^{(m,n)}_{\mathcal{SB}(B\rightarrow A)}$
@@
${\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}\subset{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}}$.
As a consequence of Proposition 3, for any $\phi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$ and $\psi\in{\mathcal{GC}}^{(m+n)}_{\mathcal{SB}}$, we have
$\phi\circ\psi\in{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}\cap{\mathcal{GC}}^{(m+n)}_{\mathcal{SB}}$,
Next, take an $(m+n)$-mode constant channel $\Theta=\Theta(0,\Gamma_{0},\mathbf{d}_{0})$. Obviously, by Theorem 7, $\Theta$ is steering-breaking. However, if $\Gamma_{0}+(0_{2m}\oplus i\Omega_{n})\ngeq 0$, and $\Gamma_{0}+(i\Omega_{m}\oplus 0_{2n})\ngeq 0$, $\Theta$ is neither steering-annihilating nor maximal Gaussian unsteerable. So
${\mathcal{GC}}^{(m+n)}_{\mathcal{SB}}\not\subset{\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}$
@@
${\mathcal{GC}}^{(m+n)}_{\mathcal{SB}}\not\subset{\mathcal{GC}}^{(m,n)}_{\mathcal{MUS}}$.
@@
${\mathcal{GC}}^{(m,n)}_{\mathcal{SA}}\not\subset{\mathcal{GC}}^{(m+n)}_{\mathcal{SB}}$.
For example, take $\tilde{\phi}=\phi\otimes I_{B}\in{\mathcal{GC}}^{(1+1)}$, where $\phi=\phi(\cos\theta I_{2},\sin^{2}\theta I_{2},\mathbf{d}_{0})$ is a single-mode attenuator channel. Then $\tilde{\phi}=\tilde{\phi}(K,M,\mathbf{d})$ can be represented by
@@
and
$M+i{\Omega_{2}}=M+i\left(\Omega_{1}\oplus\Omega_{1}\right)$
$=$ $\left(\begin{array}[]{cccc}\sin^{2}\theta&i&0&0\\ -i&\sin^{2}\theta&0&0\\ 0&0&0&i\\ 0&0&-i&0\end{array}\right)\ngeq 0.$